# DEEPTalk: Dynamic Emotion Embedding for Probabilistic Speech-Driven 3D Face Animation

### v2

@@ Abstract
Speech-driven 3D facial animation has garnered lots of attention thanks to its broad range of applications. Despite recent advancements in achieving realistic lip motion, current methods fail to capture the nuanced emotional undertones conveyed through speech and produce monotonous facial motion. These limitations result in blunt and repetitive facial animations, reducing user engagement and hindering their applicability. To address these challenges, we introduce DEEPTalk, a novel approach that generates diverse and emotionally rich 3D facial expressions directly from speech inputs. To achieve this, we first train DEE (Dynamic Emotion Embedding), which employs probabilistic contrastive learning to forge a joint emotion embedding space for both speech and facial motion. This probabilistic framework captures the uncertainty in interpreting emotions from speech and facial motion, enabling the derivation of emotion vectors from its multifaceted space. Moreover, to generate dynamic facial motion, we design TH-VQVAE (Temporally Hierarchical VQ-VAE) as an expressive and robust motion prior overcoming limitations of VAEs and VQ-VAEs. Utilizing these strong priors, we develop DEEPTalk, a talking head generator that non-autoregressively predicts codebook indices to create dynamic facial motion, incorporating a novel emotion consistency loss. Extensive experiments on various datasets demonstrate the effectiveness of our approach in creating diverse, emotionally expressive talking faces that maintain accurate lip-sync. Our project page is available at: https://whwjdqls.github.io/deeptalk.github.io/.
@@ Introduction
Speech-driven 3D facial motion generation has a wide range of applications, encompassing avatar animation for gaming or cinematic productions, virtual chatbots, and immersive virtual meetings within virtual reality environments. Despite substantial advancements in accurate lip synchronization with speech, as demonstrated by recent research (Richard et al. 2021; Fan et al. 2022; Xing et al. 2023; Stan, Haque, and Yumak 2023), the majority of these methods still produce blunt and unexpressive facial expressions. Since facial expressions are crucial for conveying nonverbal cues, this limitation significantly reduces their effectiveness in scenarios requiring realistic interactions, such as interactions with non-player characters in games or emotionally responsive virtual chatbots. Therefore, it is essential to focus on enhancing the full range of facial expressions, not just the movements of the lips.
@@
Previous studies on talking faces have generated expressions using either emotion labels (Daněček et al. 2023; Gan et al. 2023; Pan et al. 2023; Ji et al. 2021) or reference expressions (Ji et al. 2022; Tan et al. 2024; Ma et al. 2023a; Liang et al. 2022). Using emotion labels provides expressive but limited outcomes, while reference expressions offer more diversity at the cost of needing numerous expressive references. Moreover, both approaches fail to capture vocal nuances, often leading to misalignment between facial expressions and speech. As shown in Figure 1, the phrase ”That’s just what I needed!” can carry different meanings based on the emotion it conveys (e.g., happiness or anger), emphasizing the importance of facial expressions that reflect the intended sentiment. Without this alignment, expressions may appear unnaturally disconnected from the spoken words, contributing to the uncanny valley effect (Wang, Lilienfeld, and Rochat 2015).
@@
Utilizing this groundwork, we propose a talking head model, DEEPTalk, that leverages both emotion and motion priors to generate diverse emotional 3D facial expressions directly from speech inputs. We first utilize cross-modal contrastive learning to capture the emotional correlation between speech and facial expressions. Unlike (Albanie et al. 2018), which used this correlation primarily for emotion recognition from unlabeled data, we aim to use it to develop a joint embedding space, which we call DEE(Dynamic Emotion Embedding). Given the inherent ambiguity in both speech and facial expressions—where multiple expressions can correspond to a single piece of speech—we utilize probabilistic embeddings (Chun et al. 2021; Chun 2023). These embeddings are designed to model the uncertainty associated with the inputs and facilitate sampling from the probability distribution. This allows the generation of diverse emotional faces from the same speech input as illustrated by the red lines in Figure 1. After constructing a joint embedding space for facial motion and speech, we use it as a strong emotion prior to training an emotional talking head model.
@@
Building on the aforementioned robust emotion and motion priors, DEEPTalk is specifically engineered to non-autoregressively map emotional speech to our codebook indices. To ensure that generated expressions consistently reflect the input speech emotions, we introduce a novel emotion consistency loss. During training, we utilize Gumbel-Softmax (Jang, Gu, and Poole 2016) and differentiable rendering to maintain differentiability throughout the process. Our qualitative and quantitative assessments, complemented by extensive user studies, demonstrate that DEEPTalk excels at generating diverse emotional facial motions while also outperforming lip synchronization. In summary, our main contributions are as follows:
  • We design a Dynamic Emotion Embedding (DEE) by jointly learning from speech and facial motion sequences through probabilistic contrastive learning.
  • We propose a novel temporally hierarchical VQ-VAE (TH-VQVAE) to build a motion prior that is both expressive and robust.
  • We train a talking head model (DEEPTalk) using the strong priors to generate diverse and expressive emotional facial motions while outperforming SOTA models on lip synchronization.
@@
The availability of large 4D mesh datasets synchronized with audio (Richard et al. 2021; Fanelli et al. 2010) has greatly advanced deep learning-based 3D face animation, enabling robust lip synchronization. Extending these advancements, VOCA (Cudeiro et al. 2019) improves realism by generating animations from any speech inputs, employing time convolutions along with a speaker identity one-hot vector. FaceFormer (Fan et al. 2022) employs a transformer-based model to generate facial movements auto-regressively. However, despite their accurate lip synchronization, the generated upper face remains static as they are less correlated with the input speech. MeshTalk (Richard et al. 2021) addresses this issue by disentangling audio-correlated and uncorrelated movements using a categorical latent space to model upper face dynamics. Similarly, CodeTalker (Xing et al. 2023) utilizes a discrete motion prior with VQ-VAE to reconstruct dynamic facial motions across the entire face. However, these methods generate deterministic motion, overlooking the inherently non-deterministic relationship between facial motion and speech. Therefore, FaceDiffuser (Stan, Haque, and Yumak 2023) addresses the limitation of deterministic models by employing a diffusion model, which is probabilistic in nature, to generate multiple facial motions for a given speech. However, it still falls short in capturing the diverse emotional expressions conveyed through speech, which is crucial for enhancing interactiveness for talking face models.
@@
Pseudo 3D Datasets.
The scarcity of 3D datasets has led works like (Yi et al. 2023), (Yang et al. 2023), (Sung-Bin et al. 2023), EMOTE, EmoTalk, and (Fu et al. 2023) to employ SOTA reconstruction methods on diverse 2D datasets to create pseudo-3D data for body and facial motion. Due to the lack of a large-scale emotional 3D dataset, we follow this approach on a large-scale, emotionally expressive dataset MEAD. Extensive comparisons on 3D talking face datasets are provided in the Supplementary.
@@
$\hat{z}^{t}=\text{ENC}_{top}(\hat{z}^{b})\in\mathbb{R}^{\tau^{t}\times C^{t}}$ (10)
where $\tau^{b}=\frac{T}{{q^{b}}}$ and $\tau^{t}=\frac{\tau}{{q^{t}}}$ is the length of the sequence divided by bottom quant factor $q^{b}$ and length of the sequence of the bottom motion feature divided by top quant factor $q^{t}$.
Using the top motion feature, we obtain a top quantized sequence $z^{t}_{q}$ by the following
@@
Table 1: Quantitative Evaluation Results. Best performance in bold, and the second best underlined. *EmoTalk does not predict FLAME parameters preventing evaluation of FFD and was trained on RAVDESS and HDTF.
@@
We utilize MEAD for both training and testing and incorporate CREMA-D (Cao et al. 2014), RAVDESS, and HDTF (Zhang et al. 2021) for evaluation. MEAD, CREMA-D, and RAVDESS are emotional talking face video datasets recorded in lab environments with actors, while HDTF consists of talking face videos sourced from YouTube. Due to RAVDESS’s limited utterances and HDTF’s lack of emotion, we evaluate only emotion and lip sync, respectively. Additionally, MEAD’s overlapping utterance between train and test datasets make unsuitable for fair comparisons with models untrained on it, and given no clear benchmark to evaluate lip generalization on emotional in-the-wild speeches, we constructed an audio test set, Emo-Vox, derived from VoxCeleb2. To create a reliable pseudo-3D ground truth dataset, we employ the SOTA 3D face reconstruction method EMOCAv2 (Daněček, Black, and Bolkart 2022) exclusively on lab setting datasets, ensuring accurate FLAME parameter extraction. Further details on the dataset are provided in the Supplementary.
@@
We adopt FID to evaluate the realism of rendered faces. To further assess the realism of facial movements in the FLAME parameter space, we developed FFD (Frechet Face Distance), inspired by FGD (Yoon et al. 2020), and computed it on sequences of FLAME parameters using an encoder trained on MEAD and BEATv2 (Liu et al. 2024). To evaluate emotional expressiveness, we compute Emo-FID, an adaptation of FID that uses an emotion feature extractor from AffectNet (Mollahosseini, Hasani, and Mahoor 2017), replacing the inception network to focus on emotion properties. For lip-sync evaluation, we used the SyncNet score (Chung and Zisserman 2016) as Lip Vertex Error (LVE) is unsuitable for evaluating DEEPTalk’s lip generation performance. Since DEEPTalk generates diverse emotional faces, the lip vertices between the ground truth and predictions may differ even when syllables are correctly pronounced. Therefore, we employed LSE-D (Lip Sync Error Distance) and LSE-C (Lip Sync Error Confidence) from Syncnet as our evaluation metrics following (Aneja et al. 2024), as both metrics do not require ground truth lip movements. Finally, we measure Diversity (Ren et al. 2023).
@@
In Table 1, DEEPTalk achieves the highest performance on LSE-D and ranks first or second on LSE-C across the MEAD test set and other datasets, demonstrating strong generalization and precise lip synchronization. We also conducted evaluations on Emo-Vox to assess lip synchronization, while ensuring a fair comparison with both parameter-based and vertex-based models. As shown in Table 2, our model excels in LSE-D and ranks second in LSE-C, demonstrating effective generalization of lip movements to emotional, in-the-wild audio. This performance further indicates our bottom codebook’s capability to capture high-frequency details, enhancing dynamic lip movements. In contrast, despite being trained on the ground truth scan dataset, vertex-based models fail to achieve accurate lip synchronization compared to those using pseudo ground truth. Notably, FaceFormer, trained on MEAD, produces nearly neutral expressions (see Figure 3), whereas ours are more expressive, benefiting FaceFormer’s lip-sync performance due to the emotion and lip-sync trade off, as demonstrated in ablation studies.
@@
FaceDiffuser - - 21.56 10.93 0.784
Ours 1 argmax 19.23 10.65 1.085
Ours 0.1 argmax 25.48 10.60 1.156
Ours -2 argmax 25.50 10.60 1.149
Ours -4 argmax 25.55 10.60 1.155
Ours mean $\tau_{t}=1,\tau_{b}=0.1$ 9.91 10.65 1.086
Ours mean $\tau_{t}=0.1,\tau_{b}=1$ 13.03 10.65 1.091
Ours mean $\tau_{t}=4.5,\tau_{b}=1$ 23.22 10.68 1.090
Table 3: Diversity results on the MEAD test set. DEEPTalk generates diverse faces while maintaining accurate lip sync.
@@
In Figure 3, we compare our method with SOTA methods on the MEAD test set. While most methods generate natural lip movements, they sometimes misalign with the ground truth, such as opening or closing lips incorrectly. Additionally, except for EMOTE, other methods often produce incorrect or unexpressive expressions, failing to match the ground truth. This highlights the effect of DEE and $L_{emo}$ on our framework. While EMOTE generates emotional expressions using emotional labels, it often produces exaggerated faces, like unnaturally wide eyes or flattened eyebrows, which deviate from the ground truth. In contrast, DEEPTalk generates natural emotional expressions that closely resemble the ground truth faces without the need for GT emotion labels by leveraging the emotions inherent within the speech.
@@
Figure 4 illustrates the clustering of our emotion embedding by emotions, utilizing T-SNE, demonstrating that it effectively captures a meaningful emotion space. Furthermore, to evaluate its efficacy in generating emotional expressions, we conducted an experiment in which we randomly selected two audio samples with distinct emotions and interchanged their emotion embeddings from DEE’s audio encoder, before inputting them into the Face Feature Extractor. As shown in Figure 6, this swap effectively modifies the emotional expression to match the reference speech, while maintaining precise lip synchronization with the original speech. Further analysis of emotion embedding is provided in the Supplementary.
@@
We conducted A/B test with 5-point Likert scale for 98 users to evaluate model preference across two subtasks: 1) emotional alignment between speech and expression 2) lip synchronization. We sampled twenty audio clips from MEAD and ten from Emo-Vox. Detailed methodologies are provided in Supplementary. As shown in Figure 5, DEEPTalk was preferred over all other parameter-based models across all tasks except for EMOTE, which leverages ground truth emotion labels to generate faces, serving as the upper bound for emotional face generation. Note that EMOTE is deterministic and lacks expression diversity. For vertex-based models, DEEPTalk excelled in both tasks for all competitors. Due to the lack of datasets with emotional ground truth face scans and the small size of available datasets, vertex-based methods struggle with emotional expression and accurate lip synchronization. This highlights that using pseudo-3D data, as done in DEEPTalk, is a promising approach for achieving accurate emotional 3D talking face animation with precise lip synchronization.

### v3

@@ Abstract
Speech-driven 3D facial animation has garnered lots of attention thanks to its broad range of applications. Despite recent advancements in achieving realistic lip motion, current methods fail to capture the nuanced emotional undertones conveyed through speech and produce monotonous facial motion. These limitations result in blunt and repetitive facial animations, reducing user engagement and hindering their applicability. To address these challenges, we introduce DEEPTalk, a novel approach that generates diverse and emotionally rich 3D facial expressions directly from speech inputs. To achieve this, we first train DEE (Dynamic Emotion Embedding), which employs probabilistic contrastive learning to forge a joint emotion embedding space for both speech and facial motion. This probabilistic framework captures the uncertainty in interpreting emotions from speech and facial motion, enabling the derivation of emotion vectors from its multifaceted space. Moreover, to generate dynamic facial motion, we design TH-VQVAE (Temporally Hierarchical VQ-VAE) as an expressive and robust motion prior overcoming limitations of VAEs and VQ-VAEs. Utilizing these strong priors, we develop DEEPTalk, a talking head generator that non-autoregressively predicts codebook indices to create dynamic facial motion, incorporating a novel emotion consistency loss. Extensive experiments on various datasets demonstrate the effectiveness of our approach in creating diverse, emotionally expressive talking faces that maintain accurate lip-sync. Our project page is available at https://whwjdqls.github.io/deeptalk_website/
Abstract
This supplementary material includes the following sections: (1) Implementation details of our proposed models, (2) Extended analysis of our Dynamic Emotion Embedding, (3) Comparison of 3D Talking Head Datasets, (4) Details of our evaluation metrics, and (5) Additional qualitative results.
@@ Introduction
Speech-driven 3D facial motion generation has a wide range of applications, encompassing avatar animation for game or cinematic productions, virtual chatbots, and immersive virtual meetings within virtual reality environments. Despite substantial advancements in accurate lip synchronization with speech, as demonstrated by recent research (Richard et al. 2021; Fan et al. 2022; Xing et al. 2023; Stan, Haque, and Yumak 2023), most of these methods still produce blunt and unexpressive facial expressions. Since facial expressions are crucial for conveying nonverbal cues, this limitation significantly reduces their effectiveness in scenarios requiring realistic interactions, such as interactions with non-player characters in games or emotionally responsive virtual chatbots. Therefore, it is essential to focus on enhancing the full range of facial expressions, not just the lip movements.
@@
Previous studies on talking faces have generated expressions using either emotion labels (Daněček et al. 2023; Gan et al. 2023; Pan et al. 2023; Ji et al. 2021) or reference expressions (Ji et al. 2022; Tan et al. 2024; Ma et al. 2023a; Liang et al. 2022). Using emotion labels provides expressive but limited outcomes, while reference expressions offer more diversity at the cost of needing numerous expressive references. Moreover, both approaches fail to capture vocal nuances, often leading to misalignment between facial expressions and speech. As shown in Figure 1, the phrase ”That’s just what I needed!” can carry different meanings based on the emotion it conveys (e.g., happiness or anger), emphasizing the importance of facial expressions that reflect the intended sentiment. Without this alignment, expressions may appear unnatural from spoken words, contributing to uncanny valley effect (Wang, Lilienfeld, and Rochat 2015).
@@
Utilizing this groundwork, we propose a talking head model, DEEPTalk, that leverages both emotion and motion priors to generate diverse emotional 3D facial expressions directly from speech inputs. We first utilize cross-modal contrastive learning to capture the emotional correlation between speech and facial expressions. Unlike (Albanie et al. 2018), which used this correlation primarily for emotion recognition from unlabeled data, we aim to use it to develop a joint embedding space, which we call DEE(Dynamic Emotion Embedding). Given the inherent ambiguity in both speech and facial expressions—where multiple expressions can correspond to a single piece of speech—we utilize probabilistic embeddings (Chun et al. 2021; Chun 2023). These embeddings are designed to model the uncertainty associated with the inputs and facilitate sampling from the probability distribution. This allows the generation of diverse emotional faces from the same speech input as illustrated by the red lines in Figure 1. After constructing a joint embedding space for facial motion and speech, we use it as a strong emotion prior to train an emotional talking head model.
@@
Building on the aforementioned robust emotion and motion priors, DEEPTalk is specifically engineered to non-autoregressively map emotional speech to our codebook indices. To ensure that generated expressions consistently reflect the input speech emotions, we introduce a novel emotion consistency loss. Training incoporates Gumbel-Softmax (Jang, Gu, and Poole 2016) and differentiable rendering to ensure end-to-end differentiability. Our qualitative and quantitative assessments, including extensive user studies, demonstrate that DEEPTalk excels at generating diverse emotional facial motions while also outperforming lip synchronization.
In summary, our main contributions are as follows: we design a Dynamic Emotion Embedding (DEE) that jointly learns from speech and facial motion sequences through probabilistic contrastive learning. Additionally, we propose a novel temporally hierarchical VQ-VAE (TH-VQVAE) to construct a motion prior that is both expressive and robust. Finally, we train a talking head model, DEEPTalk, leveraging these strong priors to generate diverse and expressive emotional facial motions while also outperforming state-of-the-art models on lip synchronization.
@@
The availability of large 4D mesh datasets synchronized with audio (Richard et al. 2021; Fanelli et al. 2010a) has greatly advanced deep learning-based 3D face animation, enabling robust lip synchronization. Extending these advancements, VOCA (Cudeiro et al. 2019b) improves realism by generating animations from any speech inputs, employing time convolutions with a speaker identity one-hot vector. FaceFormer (Fan et al. 2022) uses a transformer-based model to generate facial movements auto-regressively. However, despite accurate lip synchronization, the generated upper face remains static as it is less correlated with input speech. MeshTalk (Richard et al. 2021) addresses this by disentangling audio-correlated and uncorrelated movements using a categorical latent space to model upper face dynamics. Similarly, CodeTalker (Xing et al. 2023) utilizes a discrete motion prior with VQ-VAE to reconstruct dynamic facial motions across the entire face. However, these methods generate deterministic motion, overlooking the inherently non-deterministic relationship between facial motion and speech. Therefore, FaceDiffuser (Stan, Haque, and Yumak 2023) addresses the limitation of deterministic models by employing a diffusion model, which is probabilistic in nature, to generate multiple facial motions for a given speech. However, it still falls short in capturing the diverse emotional expressions conveyed through speech, which is crucial for enhancing interactiveness for talking face models.
@@
where $\tau^{b}=\frac{T}{{q^{b}}}$ and $\tau^{t}=\frac{\tau}{{q^{t}}}$ is the length of the sequence divided by bottom quant factor $q^{b}$ and length of the sequence of the bottom motion feature divided by top quant factor $q^{t}$. Using the top motion feature, we obtain a top quantized sequence $z^{t}_{q}$, then use this to get decoded top features ${z}^{t}_{d}$ as
@@
Table 1: Quantitative Evaluation Results. Best performance in bold, and the second best underlined. *EmoTalk does not predict FLAME parameters preventing evaluation of FFD and was trained on RAVDESS and HDTF. LSE-C reported in Supplementary.
@@
We utilize MEAD for train and test, and incorporate CREMA-D (Cao et al. 2014), RAVDESS, and HDTF (Zhang et al. 2021) for evaluation. MEAD, CREMA-D, and RAVDESS are lab-recorded emotional talking face videos, while HDTF comprises YouTube-sourced talking face videos. Due to RAVDESS’s limited utterances and HDTF’s lack of emotion, we evaluate only emotion and lip sync, respectively. Additionally, MEAD’s overlapping utterance between train and test datasets make unsuitable for fair comparisons with models untrained on it, and given no clear benchmark for lip generalization on emotional in-the-wild speeches, we constructed an audio test set, Emo-Vox, derived from VoxCeleb2. To create a reliable pseudo-3D ground truth dataset, we employ the SOTA 3D face reconstruction method (Daněček, Black, and Bolkart 2022) exclusively on lab setting datasets for accurate FLAME parameter extraction. Further details are in the Supplementary.
@@
We adopt FID to evaluate the realism of rendered faces. To further assess the realism of facial movements in the FLAME parameter space, we developed FFD (Frechet Face Distance), inspired by FGD (Yoon et al. 2020), and computed it on sequences of FLAME parameters using an encoder trained on MEAD and BEATv2 (Liu et al. 2024). To evaluate emotional expressiveness, we compute Emo-FID, an adaptation of FID that uses an emotion feature extractor from AffectNet (Mollahosseini, Hasani, and Mahoor 2017), replacing the inception network to focus on emotion properties. For lip-sync evaluation, we used SyncNet metrics (Chung and Zisserman 2016) LSE-D (Lip Sync Error Distance) and LSE-C (Lip Sync Error Confidence) following (Aneja et al. 2024). Unlike Lip Vertex Error (LVE), which is unsuitable for DEEPTalk due to its diverse emotional faces, these metrics evaluate lip-sync without requiring ground truth lip movements. Finally, we measure Diversity (Ren et al. 2023).
@@
In Table 1, DEEPTalk achieves the highest performance on LSE-D and ranks first or second on LSE-C across the MEAD test set and other datasets, demonstrating strong generalization and precise lip sync. We also conducted evaluations on Emo-Vox to assess lip sync, while ensuring a fair comparison with both parameter-based and vertex-based models. As shown in Table 2, our model excels in LSE-D and ranks second in LSE-C, demonstrating effective generalization of lip movements to emotional, in-the-wild audio. This performance further indicates our bottom codebook’s capability to capture high-frequency details, enhancing dynamic lip movements. Moreover, due to the limited scale of the ground truth scan dataset, vertex-based models fail to achieve accurate lip sync compared to those using pseudo ground truth. Notably, FaceFormer, trained on MEAD, produces nearly neutral expressions (see Figure 3), whereas ours are more expressive, benefiting FaceFormer’s lip-sync performance due to the emotion and lip-sync trade off, as demonstrated in ablation studies.
@@
FaceDiffuser - - 21.56 10.93
Ours 1 argmax 19.23 10.65
Ours 0.1 argmax 25.48 10.60
Ours -2 argmax 25.50 10.60
Ours -4 argmax 25.55 10.60
Ours mean $\tau_{t}=1,\tau_{b}=0.1$ 9.91 10.65
Ours mean $\tau_{t}=0.1,\tau_{b}=1$ 13.03 10.65
Ours mean $\tau_{t}=4.5,\tau_{b}=1$ 23.22 10.68
Table 3: Diversity results on the MEAD test set. DEEPTalk generates diverse faces while maintaining accurate lip sync. LSE-C reported in the Supplementary.
@@
Figure 3 compares our method with SOTA methods on the MEAD test set. While most methods generate natural lip movements, they often misalign with the ground truth, such as opening or closing lips incorrectly. Additionally, except for EMOTE, other methods produce incorrect or unexpressive expressions. This highlights the effect of DEE and $L_{emo}$ on our framework. While EMOTE generates emotional expressions using emotional labels, it often produces exaggerated faces, like unnaturally wide eyes or flat eyebrows, deviating from the ground truth. In contrast, DEEPTalk generates natural emotional expressions that closely resemble the ground truth without emotion labels by leveraging emotions inherent within the speech.
@@
Figure 4 shows that our emotion embedding clusters by emotions using T-SNE, capturing a meaningful emotion space. Furthermore, to evaluate its efficacy in generating emotional expressions, we randomly selected two audio samples with distinct emotions and interchanged their emotion embeddings from DEE’s audio encoder, before inputting them into the Face Feature Extractor. As shown in Figure 6, this swap effectively modifies the emotional expression to match the reference speech, while maintaining precise lip synchronization with the original speech. Further analysis is provided in Supplementary.
@@
We conducted A/B test with 5-point Likert scale for 98 users to evaluate model preference across two subtasks: 1) emotional alignment between speech and expression 2) lip synchronization. We sampled twenty audio clips from MEAD and ten from Emo-Vox. Details are provided in Supplementary. As shown in Figure 5, DEEPTalk outperformed all parameter-based models across all tasks except for EMOTE, which leverages ground truth emotion labels to generate faces, serving as the upper bound for emotional face generation. Note that EMOTE is deterministic and lacks expression diversity. For vertex-based models, DEEPTalk excelled in both tasks for all competitors. Due to the scarcity and limited size of emotional ground truth face scan datasets, vertex-based methods struggle with emotional expression and accurate lip synchronization. This highlights that using pseudo-3D data, as done in DEEPTalk, is a promising approach for achieving accurate emotional 3D talking face animation with precise lip synchronization.